# Controllable Traffic Simulation through LLM-Guided Hierarchical Chain-of-Thought Reasoning

### v1

@@ Abstract
Evaluating autonomous driving systems in complex and diverse traffic scenarios through controllable simulation is essential to ensure their safety and reliability. However, existing traffic simulation methods face challenges in their controllability. To address this, this paper proposes a novel diffusion-based and LLM-enhanced traffic simulation framework. Our approach incorporates a unique chain-of-thought (CoT) mechanism, which systematically examines the hierarchical structure of traffic elements and guides LLMs to thoroughly analyze traffic scenario descriptions step by step, enhancing their understanding of complex situations. Furthermore, we propose a Frenet-frame-based cost function framework that provides LLMs with geometrically meaningful quantities, improving their grasp of spatial relationships in a scenario and enabling more accurate cost function generation. Experiments on the Waymo Open Motion Dataset (WOMD) demonstrate that our method handles more intricate descriptions, generates a broader range of scenarios in a controllable manner, and outperforms existing diffusion-based methods in terms of efficiency.
@@ I Introduction
To address this problem, recent work has focused on utilizing generative models to achieve controllable traffic simulation. One approach [8, 9, 10] involves using condition tokens to guide the model with basic conditions, such as goal points or velocity. However, this method requires additional training with the condition tokens and struggles to handle more complex conditions. Another approach leverages conditional sampling in diffusion models, which does not require extra training and offers more flexibility by optimizing cost functions during the sampling process. The cost functions can be either human-designed [11, 12, 13, 14] or generated by large language models (LLMs) [15]. Human-designed cost functions typically include predefined rules, such as speed limits or avoiding collisions, but are limited in their ability to address conditions outside of predefined functions. On the other hand, methods based on LLM-generated cost functions mark a significant progression to deal with a wide range of conditions. They take user descriptions of a scenario as input and create corresponding cost functions by LLMs, which is more adaptable than predefined functions. However, current LLMs still face challenges in understanding and processing intricate descriptions. Additionally, due to the complex denoiser architecture and costly sampling process, the diffusion-based approach is notably time-consuming.
Our goal is to overcome the limitations of simulation methods based on diffusion models and LLM-generated cost functions, further utilizing the powerful generative capabilities of LLMs and leveraging the flexibility of diffusion conditional sampling. To achieve this, we first propose a logical framework to guide LLMs in processing complex conditions. Drawing inspiration from research on human behavior, which highlights that humans often break actions down into subtasks [16], we introduce a novel chain-of-thought (CoT) mechanism [17]. This mechanism enables LLMs to systematically break down these events into hierarchical structures, and generate corresponding cost functions step-by-step. By systematically exploring traffic scene elements, we demonstrate the significant advantages of employing CoT to enhance understanding and analysis within this domain.
Furthermore, a proper form of cost functions is needed. Current methods project the trajectories on the surrounding lanes and utilize the projection points to build the cost functions. However, the geometrical implication of the projection is difficult for LLMs to understand, and the process of converting projection points into cost functions is also challenging for them. To address this, we propose a Frenet-frame-based cost function framework, which uses lateral and longitudinal information to build cost functions. The Frenet-based approach has clear geometric meanings and we demonstrate it can further enhance the LLM’s understanding of the scenario.
Finally, a more efficient architecture is required. Utilizing the recent techniques in motion prediction [18, 19, 20], we construct a novel traffic simulation framework. We incorporate the agent’s historical trajectories as one of the conditions, eliminating the need for complex historical data encoders. Furthermore, we use a relative coordinate system, thus our method avoids the need to re-encode invariant information like map topology during the rollout of the simulation.
@@
We evaluate our method on the Waymo Open Motion Dataset (WOMD) [21]. Our experiments demonstrate that the proposed CoT and Frenet-based functions can handle a broader range of conditions compared to existing methods, including typical driving behaviors like cut-ins, as well as abnormal behaviors such as going off-road. Experiments also show that our approach achieves faster generation speeds than the current diffusion-based methods while maintaining competitive performance in the generated scenarios.
@@
  • We present an innovative CoT mechanism designed to enhance the capability of LLMs to effectively process and interpret the hierarchical structure of elements within traffic simulations, and provide precise guidance for our generative model.
  • We develop an efficient diffusion-based traffic simulation framework with faster generation speeds than the current diffusion-based methods while maintaining competitive performance.
  • We propose a novel Frenet-frame-based cost function framework that provides clear geometric meanings, making it more understandable and usable by LLMs for building accurate cost functions.
  • Thorough experiments show that our method can process more complex descriptions and generate a broader range of scenarios in a controllable manner.
@@ II-A Traffic simulation
Diffusion-based methods. More recent work leverages diffusion models for controllable simulation, benefiting from strong generative capabilities and test-time adaptability. Most of these models [11, 12, 13, 14, 27] rely on pre-defined functions, such as goal point specification and collision avoidance, to guide the sampling process. While effective, their ability to control conditions beyond these predefined functions is constrained. CTG++ [15] marks a significant progression by utilizing large language models (LLMs) to translate user-specified descriptions into cost functions. This innovation expands the range of controllable conditions, addressing many limitations of previous approaches. However, it still faces challenges in understanding and processing complex descriptions.
@@ II-B LLMs for robotics
The application of LLMs in robotics has garnered significant attention due to their ability to bridge the gap between natural language instructions and robotic control. Recent approaches leverage LLMs to generate both actions [28, 29] and reward functions [30, 31], achieving notable progress in this area. Building on these advancements, CTG++ [15] explored the use of LLMs in traffic simulation, employing them to generate cost functions based on natural language descriptions of traffic scenarios. In this work, we aim to extend the generative capabilities of LLMs to construct more comprehensive and precise cost functions, enhancing the controllability of traffic simulations.
@@ III Preliminary
Figure 1: We propose to leverage the diffusion model to generate future trajectories in a diffusion policy manner. We further carefully design CoT-based prompting to guide the LLM to comprehend and generate the hierarchical structure of traffic scenes. Finally, the guidance generated by LLM is incorporated into the denoiser for controllable generation.
@@ III-A Problem formulation
A traffic scenario can be defined as a combination of road topology $M$ and agent states $A$. We use a set of $N_{m}$ polylines for road topology to describe lanes [26], characterized by their coordinates and attributes (such as lane type). In addition, we denote the connectivity of the polylines (predecessor, successor, left neighbor, and right neighbor) as graph edges $E=\{e_{ij}\}$, to represent the road topology. For agent states, we focus primarily on the position and heading of each agent over a sequence of timesteps. An agent’s state history and future are denoted as $A_{i}=\left\{\boldsymbol{s}_{-T_{\text{hist}}},...,\boldsymbol{s}_{0},...,\boldsymbol{s}_{T_{\text{future}}}\right\}$, where $\boldsymbol{s}_{t}=(x_{t},y_{t},\theta_{t})$. Here, $T_{\text{hist}}$ and $T_{\text{future}}$ represent the lengths of the history and future trajectories, respectively.
Our goal is to simulate $N_{a}$ agent future trajectories given (1) the road topology $M$, (2) the historical states of $N_{a}$ agents, and (3) the language description specified by the user. These generated trajectories must be consistent with the scenario’s language description.
@@ III-B Diffusion Models
Diffusion models [32] have emerged as a powerful framework to represent complex data distribution. In general, diffusion models parameterize data distribution as $p_{\theta}$ and learn to model it through score-matching, iteratively transforming a simple distribution such as Gaussian noise into a more complex data distribution through multiple small and reversible steps. The process can be modeled as stochastic differential equations (SDEs) [33] with the forward and reverse process:
$d\boldsymbol{x}=f(\boldsymbol{x},t)dt+g(t)d\omega$ (1)
$d\boldsymbol{x}=[f(\boldsymbol{x},t)-g^{2}(t)\nabla_{\boldsymbol{x}}{\log{p_{t}}(\boldsymbol{x})}]dt+g(t)d\omega$ (2)
where $\omega$ is the standard Wiener process. $\nabla_{x}{\log{p_{t}}(x)}$ is the time-dependent score, which can be approximated by training a neural network by score-matching. In our case, diffusion models are used to represent the distribution of trajectory features $x$ conditioned on map topology $p_{\theta}(x;m)$.
The core of the diffusion process includes designing a series of data distributions that are characterized by noise schedule $\sigma(t)$ and a contraction or scale schedule $s(t)$. In this work, we follow the formulation of EDM [34], which re-designed the diffusion process and depicted the sampling process by the following ODE:
$d\boldsymbol{x}=\left[\frac{\dot{s}(t)}{s(t)}\boldsymbol{x}-s(t)^{2}\dot{\sigma}(t)\sigma(t)\nabla_{\boldsymbol{x}}\log p\left(\frac{\boldsymbol{x}}{s(t)};m,\sigma(t)\right)\right]dt$ (3)
In practice, we consider the scaling schedule to be constant $s(t)=1$. Thus, the ODE in equation (3) simplifies to:
$d\boldsymbol{x}=-\dot{\sigma}(t)\sigma(t)\nabla_{x}\log p(\boldsymbol{x};m,\sigma(t))dt$ (4)
The denoiser model $D_{\theta}(\boldsymbol{x};m,\sigma)$ is trained to recover the noised sample $\boldsymbol{x}$ from its noisy version. The relationship between the score function and the denoiser model is expressed as follows:
$\nabla_{\boldsymbol{x}}\log p_{\theta}(\boldsymbol{x};m,\sigma)=\frac{D_{\theta}(\boldsymbol{x};m,\sigma)-\boldsymbol{x}}{\sigma^{2}}$ (5)
@@
$\mathbb{E}_{y\sim p_{0},\boldsymbol{n}\sim\mathcal{N}(\boldsymbol{0},\sigma^{2}\boldsymbol{I})}\left\|D_{\theta}(\boldsymbol{y}+\boldsymbol{n};m,\sigma)-\boldsymbol{y}\right\|^{2},$ (6)
where $p_{0}$ is the data distribution and $\boldsymbol{n}$ represents the Gaussian noise. After training, data can be sampled from noise using the relationship in equation (5) and the ODE described in equation (4).
@@ III-C Conditional sampling
For various problems, we have extra condition $c$ which can be described by a cost function $\mathcal{L}$. In such cases, we aim to sample data from the conditional distribution $p(\boldsymbol{x};c)$. Under our specific problem settings, this translates to sampling from $p(\boldsymbol{x};m,\sigma,c)$. Consequently, the score function should be modified from $\nabla_{\boldsymbol{x}}{\log{p}(\boldsymbol{x};m,\sigma)}$ to $\nabla_{\boldsymbol{x}}{\log{p_{t}}(\boldsymbol{x};m,\sigma,c)}$. According to Bayes’ rule, the modified score function can be expressed as:
$\nabla_{\boldsymbol{x}}{\log{p}(\boldsymbol{x};c,m,\sigma)}=\nabla_{\boldsymbol{x}}{\log{p}(\boldsymbol{x};m,\sigma)}+\nabla_{\boldsymbol{x}}{\log{p}(c;\boldsymbol{x},m,\sigma)}$ (7)
@@
$\nabla_{\boldsymbol{x}}{\log{p}(c;\boldsymbol{x},m,\sigma)}\approx\lambda\frac{\partial}{\partial\boldsymbol{x}}\mathcal{L}(D_{\theta}(\boldsymbol{x};m,\sigma)),$ (8)
where $\lambda$ is a scaling factor. By utilizing equations (4), (5) and (8), we can sample data that meets the condition $c$. In practice, to ensure the stability of the sampling process, a score threshold of $(-1,1)$ is applied to equation (8).
@@ IV Method
The overall framework of our proposed method is illustrated in Figure 1. In this section, we first introduce the Transformer-based Denoiser, which models the complex interactions within traffic scenarios (see Section IV-A). Following this, we detail the Guided Generation module. As providing clear guidance for denoiser is non-trivial, we introduce the chain-of-thought mechanism and Frenet-based cost functions (see Section IV-B) to generate the hierarchical structure of traffic scenes.
@@ IV-A Transformer-based denoiser
Relative coordinate system. In alignment with recent trajectory prediction methods [18, 19], we adopt a relative coordinate system. Initially, road polylines and trajectories are rotated and normalized to the central or current position. To further capture spatial relationships among traffic elements, the feature vector $\boldsymbol{d}_{ij}=(\Delta x_{ij},\Delta y_{ij},\cos\Delta\theta_{ij},\sin\Delta\theta_{ij})$ is extracted, where $(\Delta x_{ij},\Delta y_{ij})$ represents the relative positions within the reference frame of agent $i$, and $\Delta\theta_{ij}$ denotes the relative heading angle. Note that the feature vectors between static elements, such as road polylines, remain consistent regardless of the agents’ movements. Consequently, this approach eliminates the need to re-encode the map topology during simulation rollout, thereby enhancing inference speed.
Latent trajectory expression. Following recent studies [12], we use Principal Component Analysis (PCA) to express trajectories in the latent space. Specifically, for each trajectory $A_{i}\in\mathbb{R}^{T\times 3}$, PCA maps it into the latent representation $\tilde{A}_{i}\in\mathbb{R}^{d_{\text{PCA}}}$, where $d_{\text{PCA}}$ is the number of principal components. The latent expression captures both spatial and temporal features, enabling the model to encode trajectories directly with a Multi-Layer Perceptron (MLP) at a low computational cost, rather than employing models like attention mechanisms with high computational overhead.
History guidance. Current methods [12, 15] utilize complex encoders to model historical information, directly learning the distribution $p(x;m,A_{\text{hist}})$. Our approach, however, treats the historical states $A_{\text{hist}}$ as an additional condition rather than embedding it directly into the model. Consequently, our model learns the distribution $p(x;m)$ and employs the sampling method introduced in Section III-C to generate data conditioned on historical states. Specifically, we express this condition using the L2 norm between the generated trajectory history $\boldsymbol{y}_{\text{hist}}$ and the ground-truth history $\hat{\boldsymbol{y}}_{\text{hist}}$:
$\mathcal{L}_{\text{hist}}=\left\|\boldsymbol{y}_{\text{hist}}-\hat{\boldsymbol{y}}_{\text{hist}}\right\|^{2}$ (9)
By incorporating this additional loss term within the framework of controllable simulation, we effectively integrate historical information without the need for a separate encoder, simplifying the overall model design and enhancing computational efficiency.
@@
where $X_{m}$ is the polyline-level feature and $d$ denotes the hidden size of the model. For agent features $X_{a}\in\mathbb{R}^{N_{a}\times d}$, we utilize an MLP to encode the noise trajectories in the PCA latent space.
After feature encoding, we employ three distinct attention modules to extract the lane-lane, agent-lane, and agent-agent interactions. Let $\boldsymbol{x}_{i}$ denote the feature of an object, the attention mechanism can be expressed as:
$\tilde{\boldsymbol{x}_{i}}=$ $\sum_{j}\mathrm{Softmax}(s_{ij})(\boldsymbol{x}_{j}\boldsymbol{W}_{v}+\mathrm{MLP}(\boldsymbol{r}_{ij})),$ (11)
$s_{ij}=$ $\frac{1}{\sqrt{d}}(\boldsymbol{x}_{i}\boldsymbol{W}_{q})(\boldsymbol{x}_{j}\boldsymbol{W}_{k}+\mathrm{MLP}(\boldsymbol{r_{ij}})),$ (12)
where $\boldsymbol{W}$ represents a Linear layer. For agent-agent and agent-lane attentions, $\boldsymbol{r}_{ij}$ is the spatial relativity feature $\boldsymbol{d}_{ij}$. For lane-lane attentions, $\boldsymbol{r}_{ij}=\boldsymbol{d}_{ij}+\boldsymbol{e}_{ij}$ is the sum of the spatial relativity feature and the connectivity feature.

### v2

# Controllable Traffic Simulation through LLM-Guided Hierarchical Reasoning and Refinement

@@ Abstract
Evaluating autonomous driving systems in complex and diverse traffic scenarios through controllable simulation is essential to ensure their safety and reliability. However, existing traffic simulation methods face challenges in their controllability. To address this, we propose a novel diffusion-based and LLM-enhanced traffic simulation framework. Our approach incorporates a high-level understanding module and a low-level refinement module, which systematically examines the hierarchical structure of traffic elements, guides LLMs to thoroughly analyze traffic scenario descriptions step by step, and refines the generation by self-reflection, enhancing their understanding of complex situations. Furthermore, we propose a Frenet-frame-based cost function framework that provides LLMs with geometrically meaningful quantities, improving their grasp of spatial relationships in a scenario and enabling more accurate cost function generation. Experiments on the Waymo Open Motion Dataset (WOMD) demonstrate that our method can handle more intricate descriptions and generate a broader range of scenarios in a controllable manner.
@@ I Introduction
To address this problem, recent work has focused on utilizing generative models to achieve controllable traffic simulation. One approach [8, 9, 10] involves using condition tokens to guide the model with basic conditions, such as goal points or velocity. However, this method requires additional training with the condition tokens and struggles to handle more complex conditions. Another approach leverages conditional sampling in diffusion models, which does not require extra training and offers more flexibility by optimizing cost functions during the sampling process. The cost functions can be either human-designed [11, 12, 13, 14] or generated by large language models (LLMs) [15, 16]. Human-designed cost functions typically include predefined rules, such as speed limits or avoiding collisions, but are limited in their ability to address conditions outside of predefined functions. On the other hand, methods based on LLM-generated cost functions mark a significant progression to deal with a wide range of conditions. They take user descriptions of a scenario as input and create corresponding cost functions by LLMs, which is more adaptable than predefined functions. However, current LLMs still face challenges in either processing intricate descriptions effectively or generating outputs that are easily guidable, resulting in a low success rate for producing correct traffic simulation.
Our goal is to overcome the limitations of simulation methods based on diffusion models and LLM-generated cost functions, further utilizing the powerful generative capabilities of LLMs and leveraging the flexibility of diffusion conditional sampling. To achieve this, we propose a novel LLM-based framework consisting of two key modules: a high-level understanding module, and a low-level refinement module.
The high-level module is designed to guide LLMs in processing complex conditions and generating appropriate cost functions. Here, we systematically design a novel chain-of-thought (CoT) mechanism [17] for traffic scene analysis, which enables LLMs to systematically break down the traffic events into hierarchical structures and generate corresponding cost functions step-by-step. By reasoning the interaction graphs in the traffic scene and breaking them down into manageable components, we demonstrate the significant advantages in understanding and processing complex traffic scenes. However, the generated cost functions may occasionally be suboptimal or challenging to control via diffusion guidance. To address this, we introduce a low-level refinement module, which takes the low-level trajectory coordinates as inputs and enables the LLM to perform self-reflection and iterative improvement. This module further refines the generated cost functions, enhancing the model’s ability to comprehend complex scenes and improving the performance of guided simulation.
Moreover, the formulation of appropriate cost functions is crucial. Existing approaches often rely on target points [16] or project trajectories onto surrounding lanes and utilize the projection points [15] to construct cost functions. However, these methods present significant limitations. The use of target points overly restricts the feasible trajectory space, as a single description may correspond to multiple valid target points rather than a single one. Additionally, the geometric interpretation of projection points is inherently difficult for LLMs to comprehend, and the process of translating these projections into cost functions further complicates the task. To address these issues, we propose a novel Frenet-frame-based cost function framework. This approach leverages lateral and longitudinal information to construct cost functions, providing clear geometric interpretations and more reasonable constraints. We demonstrate that this Frenet-based method not only enhances the LLM’s understanding of the scenario but also significantly improves the success rate of diffusion guidance.
Regarding the model architecture, we integrate the state-of-the-art trajectory prediction frameworks [18, 19, 20] with a relative coordinate system to ensure translational and rotational invariance for each agent in the scene. Furthermore, we formulate the agent’s known trajectories as a cost function that serves as guidance during diffusion test-time sampling. This design provides a flexible framework that can accommodate various applications, including unconditional generation, and conditional simulation, without relying on extra training and complex models.
@@
We evaluate our method on the Waymo Open Motion Dataset (WOMD) [21]. Experimental results demonstrate that the proposed LLM framework and Frenet-based cost functions can effectively handle a wider range of driving scenarios compared to existing methods. This includes not only typical driving behaviors, such as cut-ins, but also abnormal behaviors, such as going off-road. Additionally, our method achieves zero-shot ability on accident events despite being trained solely on a nearly collision-free dataset. Extensive experiments further validate that our approach delivers competitive performance across multiple tasks.
@@
  • We propose a novel LLM-based framework designed to interpret complex traffic scenarios and generate precise cost functions, enabling more accurate and reliable traffic simulation.
  • We innovatively address the effectiveness of Frenet coordinate system in traffic simulation domain, significantly enhancing the generative capabilities of LLMs and improving the success rate of diffusion-guided scenario generation.
  • Through extensive experiments, we demonstrate that our method can handle more intricate scenario descriptions and generate a wider variety of traffic scenarios in a controllable manner, including complex interactions and accident simulations.
@@ II-A Traffic simulation
Diffusion-based methods. More recent work leverages diffusion models for controllable simulation, benefiting from strong generative capabilities and test-time adaptability. Most of these models [11, 12, 13, 14, 27] rely on pre-defined functions, such as goal point specification and collision avoidance, to guide the sampling process. While effective, their ability to control conditions beyond these predefined functions is constrained. CTG++ [15] and SceneDiffuser [16] marks a significant progression by utilizing large language models (LLMs) to translate user-specified descriptions into cost functions, which expands the range of controllable conditions. However, it still faces challenges in understanding and processing complex descriptions.
@@ II-B LLMs for robotics
The application of LLMs in robotics has garnered significant attention due to their ability to bridge the gap between natural language instructions and robotic control. Recent approaches leverage LLMs to generate both actions [28, 29] and reward functions [30, 31, 32], achieving notable progress in this area. Building on these advancements, CTG++ [15] and SceneDiffuser [16] explored the use of LLMs in traffic simulation, employing them to generate cost functions based on natural language descriptions of traffic scenarios. In this work, we aim to extend the generative capabilities of LLMs to construct more comprehensive and precise cost functions, enhancing the controllability of traffic simulations.
@@ III Preliminary
Figure 1: Method Overview. We employ LLM to generate guidance and utilize a diffusion model to produce trajectories. The high-level module generates a cost function, which guides the transformer-based diffusion model to generate trajectories. If the trajectory generation fails, the coordinates are fed into the low-level module to further refine the cost function.
@@ III-A Problem formulation
A traffic scenario can be defined as a combination of road topology $m$ and agent states $a$. We use a set of $N_{m}$ polylines for road topology to describe lanes [26], characterized by their coordinates and attributes (such as lane type). In addition, we denote the connectivity of the polylines (predecessor, successor, left neighbor, and right neighbor) as graph edges $E=\{e_{ij}\}$, to represent the road topology. For agent states, we focus primarily on the position and heading of each agent over a sequence of timesteps. An agent’s state history and future are denoted as $a_{i}=\left\{\bm{s}_{-T_{\text{hist}}},...,\bm{s}_{0},...,\bm{s}_{T_{\text{future}}}\right\}$, where $\bm{s}_{t}=(x_{t},y_{t},\theta_{t})$. Here, $T_{\text{hist}}$ and $T_{\text{future}}$ represent the lengths of the history and future trajectories, respectively.
Our goal is to simulate $N_{a}$ agent future trajectories given (1) the road topology $m$, (2) the current or historical states of $N_{a}$ agents, and (3) the language description specified by the user. These generated trajectories must be consistent with the scenario’s language description.
@@ III-B Diffusion Models
Diffusion models [33] have emerged as a powerful framework to represent complex data distribution. In general, diffusion models parameterize data distribution as $p_{\theta}$ and learn to model it through score-matching, transforming a simple distribution (e.g., Gaussian noise) into a complex data distribution via multiple small, reversible steps. This process is modeled by stochastic differential equations (SDEs) [34]:
$d\bm{x}=f(\bm{x},t)dt+g(t)d\omega$ (1)
$d\bm{x}=[f(\bm{x},t)-g^{2}(t)\nabla_{\bm{x}}{\log{p_{t}}(\bm{x})}]dt+g(t)d\omega$ (2)
where $\omega$ is the standard Wiener process. $\nabla_{x}{\log{p_{t}}(x)}$ is the time-dependent score, which can be approximated by a neural network. In our case, diffusion models are used to represent the distribution of trajectory features $x$ conditioned on map topology $p_{\theta}(x;m)$.
Following EDM [35], the sampling process in this work is described by the ODE:
$d\bm{x}=-\dot{\sigma}(t)\sigma(t)\nabla_{x}\log p(\bm{x};m,\sigma(t))dt$ (3)
The denoiser model $D_{\theta}(\bm{x};m,\sigma)$ is trained to recover the noised sample $\bm{x}$ from its noisy version. The relationship between the score function and the denoiser model is expressed as follows:
$\nabla_{\bm{x}}\log p_{\theta}(\bm{x};m,\sigma)=\frac{D_{\theta}(\bm{x};m,\sigma)-\bm{x}}{\sigma^{2}}$ (4)
@@
$\mathbb{E}_{y\sim p_{0},\bm{n}\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I})}\left\|D_{\theta}(\bm{y}+\bm{n};m,\sigma)-\bm{y}\right\|^{2},$ (5)
where $p_{0}$ is the data distribution and $\bm{n}$ represents the Gaussian noise. After training, data can be sampled from noise using the relationship in equation (4) and the ODE described in equation (3).
@@ III-C Conditional sampling
For various problems, we have extra condition $c$ which can be described by a cost function $\mathcal{L}$. In such cases, we aim to sample data from the conditional distribution $p(\bm{x};c)$. Under our specific problem settings, this translates to sampling from $p(\bm{x};m,\sigma,c)$. Consequently, the score function should be modified from $\nabla_{\bm{x}}{\log{p}(\bm{x};m,\sigma)}$ to $\nabla_{\bm{x}}{\log{p_{t}}(\bm{x};m,\sigma,c)}$. According to Bayes’ rule, the modified score function $\nabla_{\bm{x}}{\log{p}(\bm{x};c,m,\sigma)}$ can be expressed as:
$\nabla_{\bm{x}}{\log{p}(\bm{x};m,\sigma)}+\nabla_{\bm{x}}{\log{p}(c;\bm{x},m,\sigma)}$ (6)
@@
$\nabla_{\bm{x}}{\log{p}(c;\bm{x},m,\sigma)}\approx\lambda\frac{\partial}{\partial\bm{x}}\mathcal{L}(D_{\theta}(\bm{x};m,\sigma)),$ (7)
where $\lambda$ is a scaling factor. By utilizing equations (3), (4) and (7), we can sample data that meets the condition $c$. In practice, to ensure the stability of the sampling process, a score threshold of (1,1)(-1,1)( - 1 , 1 ) is applied to equation (7).
@@ IV Method
The overall framework of our proposed method is illustrated in Figure 1. In this section, we first introduce the Transformer-based Denoiser, which models the complex interactions within traffic scenarios (see Section IV-A). Following this, we detail the LLM-based Guided Generation module in Section IV-B.
@@ IV-A Transformer-based denoiser
Relative coordinate system. In alignment with recent trajectory prediction methods [18, 36], we adopt a relative coordinate system. Initially, road polylines and trajectories are rotated and normalized to the central or current position. To further capture spatial relationships among traffic elements, the feature vector $\bm{d}_{ij}=(\Delta x_{ij},\Delta y_{ij},\cos\Delta\theta_{ij},\sin\Delta\theta_{ij})$ is extracted, where $(\Delta x_{ij},\Delta y_{ij})$ represents the relative positions within the reference frame of agent $i$, and $\Delta\theta_{ij}$ denotes the relative heading angle. Note that the feature vectors between static elements, such as road polylines, remain consistent regardless of the agents’ movements. Consequently, this approach eliminates the need to re-encode the map topology during simulation rollout, thereby enhancing inference speed.
Trajectory conditional guidance. Our approach treats the known trajectory states $A_{\text{known}}$ as an additional condition rather than embedding it directly into the model. Consequently, our model learns the distribution $p(x;m)$ and employs the sampling method introduced in Section III-C to generate data conditioned on historical states. Let $\mathcal{T}=\{t_{1},t_{2},\dots,t_{k}\}$ denote the set of timesteps where condition states are provided. The predicted trajectory states at these timesteps are represented as $\bm{y}_{\text{known}}=\{\bm{y}_{t_{i}}\mid t_{i}\in\mathcal{T}\}$ , and the corresponding ground-truth states are $\hat{\bm{y}}_{\text{known}}=\{\hat{\bm{y}}_{t_{i}}\mid t_{i}\in\mathcal{T}\}$. The condition is expressed using the L2 norm:
$\mathcal{L}_{\text{known}}=\left\|\bm{y}_{\text{known}}-\hat{\bm{y}}_{\text{known}}\right\|^{2}$ (8)
By incorporating this loss term, we effectively integrate trajectory conditions without requiring separate encoders and additional training. This design enables our method to flexibly handle various tasks, such as unconditional prediction and conditional simulation.
@@
where $X_{m}$ is the polyline-level feature and $d$ denotes the hidden size of the model. For agent features $X_{a}\in\mathbb{R}^{N_{a}\times d}$, we embed the trajectories into PCA space following [12] and then utilize an MLP to encode the noise trajectories.
After feature encoding, we employ three distinct attention modules to extract the lane-lane, agent-lane, and agent-agent interactions. Let $\bm{x}_{i}$ denote the feature of an object, the attention mechanism can be expressed as:
$\tilde{\bm{x}_{i}}=$ $\sum_{j}\mathrm{Softmax}(s_{ij})(\bm{x}_{j}\bm{W}_{v}+\mathrm{MLP}(\bm{r}_{ij})),$ (10)
$s_{ij}=$ $\frac{1}{\sqrt{d}}(\bm{x}_{i}\bm{W}_{q})(\bm{x}_{j}\bm{W}_{k}+\mathrm{MLP}(\bm{r_{ij}})),$ (11)
where $\bm{W}$ represents a Linear layer. For agent-agent and agent-lane attentions, $\bm{r}_{ij}$ is the spatial relativity feature $\bm{d}_{ij}$. For lane-lane attentions, $\bm{r}_{ij}=\bm{d}_{ij}+\bm{e}_{ij}$ is the sum of the spatial relativity feature and the connectivity feature.